# Enabling Intelligent Interactions between an Agent and an LLM: A Reinforcement Learning Approach

### v6

@@ 5.2.1 Can our agent complete target tasks with less interaction costs?
We compare our proposed approach When2Ask with several baseline methods to evaluate its effectiveness. We analyze the learning curves for both communication costs (Fig. 5) and task performances (Fig. 6) across all five environments. Additionally, we provide asymptotic performances in Table 1. As is shown, our approach successfully reduces the number of interactions with the LLM while maintaining task performance across all environments. This reduction in communication cost indicates that our method effectively learns to minimize non-informative interactions with the LLM. Furthermore, our approach maintains consistently high success rates throughout the learning process. This observation indicates that the asking policy learns to filter out unnecessary interactions while still engaging in essential interactions with the LLM to achieve successful task completion.
Figure 5: The number of interactions with the LLM vs. the number of RL iterations used for learning the asking policy. It shows that, for every environment, the more thoroughly the asking policy is trained, the fewer interactions with the LLM planner (i.e., the less interaction costs) are required to complete task.
@@ 5.2.2 Can our agent proactively seek assistance from an LLM in exploratory environments?
Upon analyzing the agent’s performance in situations where it is expected to ask the LLM planner for help, we observe that the baseline method with a hard-coded asking policy exhibited significantly lower success rates compared to other approaches. This discrepancy occurs because the agent continues executing every option until its termination condition is met, even when it has already gathered sufficient information to complete the task. Consequently, this inefficient approach results in wasted time on each option and ultimately leads to failure in completing the task within the given time limit. In contrast, our proposed approach, along with other baseline methods, demonstrates the ability to early-stop options when necessary. As a result, they achieve 100 percent success rates in SimpleDoorKey and KeyInBox.
In a specific scenario within the ColoredDoorKey environment, as illustrated in Fig. 7, we see an interesting phenomenon. The agent has chosen to take the Explore option and acquired information about the location of the yellow key (frame 2). With use of the Hard-coded baseline approach, the agent shall continue with the Explore option until it has fully explored the entire room. In contrast, using our proposed approach, the agent can recognize the value of asking the LLM planner for guidance given the current information, and immediately propose asking about the next steps while ceasing further exploration. The LLM would instruct the agent to efficiently pick up the yellow key without wasting additional time. This example highlights the effectiveness of our proposed approach in recognizing when to seek assistance from the LLM planner and making more efficient decisions based on the available information. By leveraging the embedded knowledge of the LLM planner, our approach enables the agent to make informed choices that expedite task completion and improve overall performance.
[…]
@@
5.2.3 How does the performance of our agent compare to a baseline RL agent that does not utilize an LLM planner?
To assess the importance of the reasoning ability of the LLM in our approach, we conduct an ablation study comparing it with a baseline RL agent that does not use an LLM planner. The RL baseline focuses on learning the planner, specifically policy-over-options, without any interaction with an LLM. The summarized results displayed in Table 2 demonstrate that even in the simplest environment, SimpleDoorKey, the RL baseline faces challenges in completing the task within a fixed number of training iterations. This suggests that learning how to solve these tasks from scratch is difficult for an RL agent. In embodied environments like these, agents must acquire skills such as exploration, reasoning about relationships between objects, and planning optimal actions to accomplish tasks successfully. By incorporating the LLM’s assistance, an agent can leverage the world knowledge embedded in the language model, leading to a significant reduction in the difficulties associated with solving these tasks. Consequently, the outcomes of the ablation study support the notion that the reasoning ability provided by the pre-trained LLM plays a crucial role in achieving higher performance in complex environments.
Table 2: Performance comparison between our agent and an RL agent that does not use LLM in the SimpleDoorKey environment.
Performance metric RL Our approach
Average Return $\uparrow$ 0.0324 0.7583
Average # of state-transition timesteps $\downarrow$ 98.36 30.47
Success rate $\uparrow$ 12% 100%
@@ 5.3 Habitat Experiments
Stage1 success ratenormal-$\uparrow$ 10.8% 7.6% 53.6%
Stage2 success ratenormal-$\uparrow$ 2.4% 1.6% 46.4%
Stage3 success ratenormal-$\uparrow$ 2.0% 1.2% 35.6%
Total # of interactionsnormal-$\downarrow$ 1.00 295.60 7.99
@@
We compare our approach against baselines on the Pick&Place task. To ensure reliability of experimental results, we utilize 10 training seeds to initialize the policy network. This allows us to explore different initializations and avoid biased results. Subsequently, we select the best policy obtained from these training runs to run 250 independent testing trials. As presented in Table 3 and Fig. 8, our approach significantly outperforms baseline methods across all three stages. Particularly, compared to the hard-coded baseline where the preset plan is executed step-by-step, our approach addresses the “hand-off problem" Szot et al. (2021) that can arise when the preceding option terminates at a state that makes it challenging for the succeeding option to initiate. This issue is depicted in Fig. 9, where the robot stops at an unfavorable location at the end of the Navigate option, resulting in a failure to execute the subsequent Pick option. Our approach effectively bypasses this problem by incorporating intelligent interactions with the LLM planner, enabling the agent to adapt its actions based on dynamic information provided by the planner.
@@
The obtained results demonstrate that the RL learned asking policy effectively establishes a connection between the world knowledge embedded within the LLMs and the local knowledge embedded within the pre-trained skills. This connection leads to a superior overall performance of our approach compared to the baselines that do not involve any learning. These findings align with the main observations from our experiments in the MiniGrid environments, particularly in the ColoredDoorKey scenario, where the RL learned asking policy enables the agent to outperform all baselines.
6 Conclusions
In this paper, we propose a novel RL approach that intelligently manages interactions between local agents and LLMs, taking into account both interaction cost and task completion efficiency. Our approach enables a local agent to learn when to initiate or maintain interaction with the LLM, as well as when to rely on its own learned skills without requiring LLM interaction. It offers a pathway for more efficient and intelligent use of LLMs in diverse real-world applications. We conduct extensive experiments to demonstrate the effectiveness of our approach, yielding two major insights. Firstly, RL can provide an elegant and practical solution to optimize the trade-off between interaction cost and task completion efficiency in the context of LLM-assisted sequential decision making. Secondly, by minimizing unnecessary interactions with the LLM, the agent can achieve more robust performance. This observation aligns with the learning process in humans, where excessive reliance on external forces, such as teachers or experts, can hinder the improvement of task completion abilities, especially in terms of generalizing to new tasks. Conversely, by utilizing existing abilities to handle new situations or tasks and seeking external assistance only when necessary, individuals can enhance their ability to improve at a faster pace.
@@ Acknowledgments
This work is supported by Exploratory Research Project (No.2022RC0AN02) of Zhejiang Lab.

### v8

@@ 5.2.1 Can our agent complete target tasks with less interaction costs?
We compare our approach When2Ask with baseline methods to evaluate its effectiveness. We analyze the learning curves for both interaction costs (Fig. 5) and task performances (Fig. 6) across all five environments. Additionally, we provide asymptotic performances in Table 1. As is shown, our approach successfully reduces the number of interactions with the LLM while maintaining task performance across all environments. This reduction in interaction cost indicates that our method effectively learns to reduce non-informative interactions with the LLM. Furthermore, our approach maintains consistently high success rates throughout the learning process. This observation indicates that the asking policy learns to filter out unnecessary interactions while still engaging in essential ones to achieve successful task completion.
Figure 5: The number of interactions with the LLM vs. the number of RL iterations used for learning the asking policy. It shows that, for every environment, the more thoroughly the asking policy is trained, the fewer interactions with the LLM planner (i.e., the less interaction costs) are required to complete the task. The shaded areas within the curves represent confidence intervals based on three standard errors.
@@ 5.2.2 Can our agent proactively seek assistance from an LLM in exploratory environments?
Upon analyzing the agent’s performance in situations where it is expected to ask the LLM planner for help, we observe that the baseline method with a hard-coded asking policy exhibited significantly lower success rates compared to other approaches. This discrepancy occurs because the agent continues executing every option until its termination condition is met, even when it has already gathered sufficient information to complete the task. Consequently, this inefficient approach results in wasted time on each option and ultimately leads to failure in completing the task within the given time limit. In contrast, When2Ask, along with other baseline methods, demonstrates the ability to early-stop options when necessary. As a result, they achieve 100 percent success rates in SimpleDoorKey and KeyInBox.
In a specific scenario within the ColoredDoorKey environment, as illustrated in Fig. 7(a), we see an interesting phenomenon. The agent has chosen to take the Explore option and acquired information about the location of the yellow key (frame 2). With use of the Hard-coded baseline approach, the agent shall continue with the Explore option until it has fully explored the entire room. In contrast, using our proposed approach, the agent can recognize the value of asking the LLM planner for guidance given the current information, and immediately propose querying the LLM planner for an updated plan while ceasing further exploration. The LLM would instruct the agent to efficiently pick up the yellow key without wasting additional time. This example highlights the effectiveness of our approach in recognizing when to seek assistance from the LLM planner and making more efficient decisions based on the available information. By leveraging the power of the LLM planner, our approach enables the agent to make informed choices that expedite task completion and improve overall performance.
(a) An example scenario where the agent discovers new information during option explore.
(b) An example scenario where the hard-coded translator fails to encode all information.
@@ 5.3 Habitat Experiments
Stage1 success $\uparrow$ 10.8% 7.6% 53.6%
Stage2 success $\uparrow$ 2.4% 1.6% 46.4%
Stage3 success $\uparrow$ 2.0% 1.2% 35.6%
Total # of $\downarrow$ 1.00 295.60 7.99
@@
We compare our approach against baselines on the Pick&Place task. To ensure reliability of experimental results, we utilize 10 training seeds to initialize the policy network. This allows us to explore different initializations and avoid biased results. Subsequently, we select the best policy obtained from these training runs to run 250 independent testing trials. As presented in Table 2 and Fig. 8, our approach significantly outperforms baselines across all three stages. Particularly, compared to the hard-coded baseline where the preset plan is executed step-by-step, our approach is significantly better at addressing the “hand-off problem” Szot et al. (2021) that can arise when the preceding option terminates at a state that makes it challenging for the succeeding option to initiate. This issue is depicted in Fig. 9, where the robot stops at an unfavorable location at the end of the Navigate option, resulting in a failure to execute the subsequent Pick option. Our approach effectively bypasses this problem by seeking guidance from the LLM planner.
@@
The obtained results demonstrate that the RL learned asking policy effectively establishes a connection between the world knowledge embedded within the LLMs and the downstream fine-grained knowledge embedded within the agent. This connection leads to a superior overall performance of our approach compared to the baselines that do not involve any learning. These findings align with the main observations from our experiments associated with the MiniGrid environments, particularly in the ColoredDoorKey scenario, where the RL learned asking policy enables the agent to outperform all baselines in terms of task completion success rate.
6 Concluding Remarks
We examine the application of RL in acquiring “mediator” policies for instruction-following agents powered by LLMs. Prior research has indicated that LLMs, when coupled with well-constructed prompts, can effectively generate high-level instructions conditional on state descriptions to devise detailed plans for task completion. Recent frameworks for LLM-driven planning have explored two primary strategies: 1) generating an updated plan at each timestep, and 2) requesting an update only after each plan (option) concludes based on predefined termination criteria. The former method is computationally intensive due to the expense of generating a new response from the LLM, while the latter may encounter challenges as an ongoing plan executed cannot be interrupted in time to respond to new observations.
To this end, we propose When2Ask, which involves training a mediator policy to determine when to prompt the LLM to generate an appropriate plan for the present moment. When2Ask trains the mediator policy to optimize task-oriented rewards while penalizing cases where the planner was invoked but returned the same plan the agent was already following. Experiment results across different embodied environments illustrate that the learned mediator policies achieve comparable task success rates to fixed policies that query the LLM at each timestep, while significantly reducing the number of queries to the LLM and consequently lowering the computational burden on the agent.
The utilization of LLMs to furnish robots and other autonomous agents with general-purpose reasoning and planning capabilities has shown great potential. However, this potential is somewhat limited by the quality of the mapping between low-level observations and actions, and the high-level LLM-based planner. A long-term goal would be to learn this mapping in an end-to-end way. In the interim, however, it is worth to investigate how different elements of this mapping can be learned, and how much benefit can be gained from such an endeavor, as demonstrated in this study.
@@ Acknowledgments
We appreciate anonymous reviewers for valuable comments that helped us to improve the quality of this paper. This work was primarily supported by Exploratory Research Project (No.2022RC0AN02) of Zhejiang Lab, and partly supported by China Postdoctoral Science Foundation (No.2023M743266), and Zhejiang Provincial Postdoctoral Research Project (No.ZJ2023067). Z. Xu was partly supported by Fudan University (No.JIH2325015Y).
Appendix A Appendix
A.1 Experimental settings on MiniGrid
In the basic setups of SimpleDoorKey and KeyInBox, each room contains only one key and one locked door. In SimpleDoorKey, the key is placed on the floor, while in KeyInBox, the key is inside a box. The agent needs to explore the room to locate the target door and the key/box, pick up the key, and finally use the key to unlock the target door.
In the RandomBoxKey environment, the placement of the key is randomized, either on the floor or inside a box. The agent needs to actively plan its actions based on the feedback from the environment, adjusting its plan depending on whether it observes a key or a box.
ColoredDoorKey introduces multiple keys and only one exit door. Each key and its corresponding door are color-coded, requiring a matching-colored key to unlock the door. This environment tests the agent’s ability to identify and utilize color information for successful task completion.
MovingObstacle adds another layer of complexity by introducing obstacles that move randomly within the room, potentially blocking the agent’s path. The agent needs to navigate in this dynamically changing environment and adapt its plans accordingly based on new observations.
A.1.1 More details on the design of our agent
Actor  In our experiments, the actor comprises a set of pre-defined option policies. The available options are as follows:
  • Explore: This option is implemented using preset rules. The specific procedure is as follows: first, the agent is instructed to move to the top-left corner of the task environment, and then proceed to traverse each row alternately while scanning, until all information within the environment becomes visible. This option enables the agent to uncover unexplored areas for discovering new information.
  • Go to [an object]: With this option, the agent can navigate to an object within the environment. The object can be any interactable element, such as a key, box, or door.
  • Pickup [an object]: This option enables the agent to pick up a specified object. It is useful when the agent needs to acquire an item to progress in the task, like grabbing a key to unlock a door.
  • Toggle [an object]: Using this option, the agent can the state of a particular object. Examples include opening or closing a door, use a key to unlock a door or open a box.
These pre-defined options provide the agent with a repertoire of high-level actions to choose from during its decision-making process. By selecting the appropriate option based on its current objective and observations, the agent can efficiently navigate and interact with the environment to accomplish the given task. For more details, refer to the supplement materials.
[…]
A.2 Our approach can tolerate the imperfection of the translator in the mediator module
In the complex environment of ColoredDoorKey in MiniGrid, the baseline interaction method Always has been observed to fail in certain corner cases due to flaws of other components within the framework. Fig. 7(b) presents an example scenario in ColoredDoorKey that showcases such a case. In the first frame, the agent is instructed to go to then pick up the key. After taking a left turn to drop the carried purple key (frame 2), the LLM instructs the agent once again with go to then pick up the key, where the agent should proceed to pick up the yellow key. However, the Always baseline fails in this case because the translator does not encode information about the relative position between the agent and the target object accurately. Consequently, the translator returns the same observation [observed yellow key, observed yellow door, carrying purple key] for both frames 1 and 2. In contrast, our approach learns “not to ask" for assistance in this particular case, allowing the agent to complete the action of picking up the yellow key before requesting further instructions. This highlights a significant advantage of our approach over baseline methods: it can adapt to situations where the translator’s translation process loses a lot of information. The learned asking policy enables the agent to make more informed decisions based on its observations and context, leading to robust performance in scenarios where baseline methods may fail due to flaws of the translator.
A.3 Experimental setting on Habitat
Habitat is a simulation platform specifically designed for end-to-end development of embodied AI Szot et al. (2021). It provides a framework for defining various embodied AI tasks such as navigation, object rearrangement, and question answering. Additionally, it allows for the configuration of embodied agents with specific physical forms and sensors. Agents can be trained using either imitation or reinforcement learning techniques. In our experiments, we demonstrate that our approach can generalize effectively to visually realistic domains by conducting experiments within the Habitat environment.
In our experiments, we focus on the manipulation task known as Pick&Place. In this task, the robot agent’s objective is to pick up an object from a desk and precisely place it into a designated target receptacle, which in this case is the kitchen sink. The task setting is the same as in the Habitat experiment in Zhou et al. (2024).
The robot agent is equipped with a wheeled base, a 7-degree-of-freedom (DoF) arm manipulator, and a parallel-jaw gripper. Additionally, it features a camera mounted on its “head" that provides a field of view of $90^{\circ}$ and captures visual data at a resolution of $256\times 256$ pixels. As a result, the observation space of the environment comprises a visual observation denoted as $o_{v}\in\mathbb{R}^{256\times 256\times 1}$ from the depth camera. It also includes a sensor observation $o_{s}\in\mathbb{R}^{24}$ sourced from various sensors such as joint sensors, gripping sensors, the end effector of the arm, object and target GPS sensors, among others. The action space in our setup is 11-dimensional, consisting of 3 actions controlling the robot positions, 7 actions controlling the robot arm and one action indicating termination. This action space enables the agent to execute precise movements and manipulations necessary for accomplishing the target task.
To effectively train each option, we design the rewards based on rearrangement measures. These measures take into account various factors such as the force exerted by the articulated agent, the distance between the object and the goal, and the angle between the agent and the goal. The specific details of these measures can be found in the Habitat documentations Szot et al. (2021).
In the Pick&Place environment, as solving the task requires progressively achieving several sub-goals, we use a composite stage reward system. More specifically, picking up the object successfully is referred to as Stage1 Completion and rewards a value of 1. Achieving navigation to the sink with the object is referred to as Stage2 Completion and also rewards a value of 1. Finally, successfully placing the apple into the target sink is referred to as Stage3 Completion and grants a higher reward value of 5. It is important to note that if any of the high-level options exceed their designated time limit, the task may terminate prematurely.
A.3.1 Implementation details of our approach on Habitat
Planner  We employ the pre-trained Vicuna-7b model as the LLM planner in our approach. In terms of prompt design, we begin by furnishing a concise instruction that conveys information about the target task. Subsequently, we provide a description of the current observation in the form of a Python list. An example of the dialogue generated by the LLM planner can be found in Fig. 10.
Figure 10: An example of the prompts and interactions for the habitat environment. Prefix prompt only contains a short task instruction.
Actor  In our experiments, we use three high-level options: {Navigate, Pick, Place}, each pre-trained with RL independently. Whenever there is a transition between these options, an automatic execution of the default action Reset Arm occurs. To ensure effective training of these options, we use 32 distinct training environment specifications with different object locations and target locations. Additionally, the agent’s initial positions are randomly generated each time the environment is reset, guaranteeing variability in training scenarios. For each option, we employ a ResNet18 backbone combined with a 2-layer LSTM architecture to train the corresponding models. During testing, the success rates of Navigate, Pick, and Place are 84%, 92%, and 91% respectively. These pre-trained models remain fixed throughout the task, ensuring consistency and stability during execution.
Training of the asking policy  Similar to our Minigrid experiment, we stack five consecutive frames of observations as inputs to the asking policy. This enables the network to capture temporal information and make informed decisions based on past observations. The network architecture for the asking policy consists of three CNN layers for embedding visual observations, one MLP layer for embedding sensor observations, and two additional MLP layers to output the logits for the binary question of whether to ask or not ask.